OpenStudy (anonymous):

Let f(x)=3x-6 and g(x)=x-2. Find f/g and its domain

OpenStudy (anonymous):

Can I do synthetic division on this problem?

OpenStudy (ineedhelplz):

Factor, f(x). Cancel common factor with g(x). Set g(x)=0 to find restriction on domain.

OpenStudy (anonymous):

So first I break down the 3x-6 to equal x-2?

OpenStudy (jdoe0001):

\(\bf f(x)=3x-6\qquad g(x)=x-2 \\ \quad \\ \cfrac{f(x)}{g(x)}\iff \cfrac{3x-6}{x-2} \)

OpenStudy (ineedhelplz):

f/g(x) = [3(x - 2)]/(x - 2) = 3 The domain is all real numbers except x ≠ 2

OpenStudy (anonymous):

Ohhhhh okay. Thanks y'all!